OpenStudy (anonymous):

The area of a rectangle is 42 square inches. The length of the rectangle is one inch more than the width. Find the dimensions.

TheSmartOne (thesmartone):

x(x+1) where x is width x(x+1)=42 x^2+x=42 x^2+x-42=0